Session: Jun 2007 (session 2)
Wavelength: 5 cm
To give feedback on this experiment, click
Medicina (mauro at Wed Jun 13 19:07:28 2007) Success :
some clouds in the sky
Onsala ( Michael Lindqvist at Wed Jun 13 19:12:11 2007) Success :
Bad weather, rain. Otherwise no known problems.
Hartebeesthoek (Jonathan Quick at Wed Jun 13 20:01:15 2007) Success :
No known problems
Torun (K. Borkowski at Thu Jun 14 07:00:48 2007) Minor failures :
There was a power failure at 16:51 that lasted 8 minutes. Good observations resumed in about an hour time.
Effelsberg (A. Polatidis at Thu Jun 14 14:24:52 2007) Minor failures :
Lost scans 1530-1608 due to cable unwrapping. Otherwise OK.
Noto (P. Cassaro at Mon Jun 18 13:16:20 2007) Success :
Effelsberg (dgraham at Fri Jun 22 07:38:13 2007) Failure : RECE
Suspect operator overrode the frequency from the schedule and set up for OH at 6032MHz, therefore high probability that all is lost.